OpenStudy (anonymous):

how can we express 1/2 under root 2 as a power of 2??? i am stuck!!

OpenStudy (saifoo.khan):

\[\Huge (\frac{1}{2}\sqrt{x})^2\]

OpenStudy (anonymous):

you really like writing big :)

OpenStudy (saifoo.khan):

Yesh! xD

OpenStudy (anonymous):

0.5(rootx)^2.

OpenStudy (anonymous):

how u write so much big ?

OpenStudy (anonymous):

possibly type it into Word, copy, and paste

OpenStudy (anonymous):

\[\sqrt{\frac{1}{2}}=(\frac{1}{2})^{\frac{1}{2}}\]

OpenStudy (anonymous):

\[=2^{-\frac{1}{2}}\]